### v3

@@
Abstract
Molecular docking is a pivotal process in drug discovery. While traditional techniques rely on extensive sampling and simulation governed by physical principles, these methods are often slow and costly. The advent of deep learning-based approaches has shown significant promise, offering increases in both accuracy and efficiency. Building upon the foundational work of FABind, a model designed with a focus on speed and accuracy, we present FABind+, an enhanced iteration that largely boosts the performance of its predecessor. We identify pocket prediction as a critical bottleneck in molecular docking and propose a novel methodology that significantly refines pocket prediction, thereby streamlining the docking process. Furthermore, we introduce modifications to the docking module to enhance its pose generation capabilities. In an effort to bridge the gap with conventional sampling/generative methods, we incorporate a simple yet effective sampling technique coupled with a confidence model, requiring only minor adjustments to the regression framework of FABind. Experimental results and analysis reveal that FABind+ remarkably outperforms the original FABind, achieves competitive state-of-the-art performance, and delivers insightful modeling strategies. This demonstrates FABind+ represents a substantial step forward in molecular docking and drug discovery. Our code is in this https URL.
1 Introduction
Molecular docking is a foundational technique in drug discovery Morris et al. (1996); Morris & Lim-Wilby (2008); Thomsen & Christensen (2006); AI4Science & Quantum (2023), serving as a computational strategy to predict the preferred orientation of ligands when bound to a protein target. This predictive process is crucial for the identification and optimization of compounds with therapeutic potential. Traditional methodologies Friesner et al. (2004); Trott & Olson (2010); Morris et al. (1996); Verdonk et al. (2003) heavily rely on exhaustive sampling and simulation techniques, which are rooted in the principles of physics and chemistry Shoichet & Kuntz (1993); Bernacki et al. (2005); Perez et al. (2016); Huang & Jacobson (2007). These approaches aim to mimic the complex interactions between ligands and their protein receptors to forecast optimal conformations. Despite their extensive application, the classical methods are often criticized for their computational intensity, which translates into slow processing and significant resource consumption Pagadala et al. (2017); Huang et al. (2006).
With the rise of computational power and machine learning, deep learning approaches for molecular docking have emerged Zhang et al. (2023a); Lu et al. (2023); Zhou et al. (2023). These methods leverage neural networks and advanced algorithms to analyze biochemical data. Deep learning methods are categorized as generative models that sample candidates selected by confidence models Corso et al. (2023b); Wang et al. (2023); Guo et al. (2023), similar to traditional methods, or regression approaches that directly predict coordinates Stärk et al. (2022); Zhang et al. (2022) or distance matrices followed by coordinate recovery Masters et al. (2022); Lu et al. (2022). Deep learning promises gains in docking accuracy and efficiency. Notably, FABind Pei et al. (2023) specifically accelerates docking while maintaining prediction precision.
Despite advances by FABind in rapid regression-based prediction, its accuracy was observed to be not fully aligned with the latest progress Corso et al. (2023b); Liu et al. (2023); Yan et al. (2023). Here, we introduce FABind+, which features enhanced pocket prediction capabilities and high-quality pose generation. Our first analysis highlights pocket prediction as a pivotal factor, where inaccuracies disrupt the docking process. As shown in Figure 1, the incorrect pocket from FABind leads to a large root-mean-square deviation (RMSD) score, with 16.25% of pose predicted by FABind having an RMSD greater than 10Å. Obviously, this issue can be attributed to incorrect pocket prediction. To address this, we propose dynamically predicting the pocket radius instead of using a fixed-size sphere, thereby allowing the enlarged pocket to cover the entire potential ligand structure for accurate docking pose prediction. By inheriting the pocket prediction task from FABind, we aim to create a more suitable region for docking by meticulously tuning the multitask framework. For the docking module, we also implement strategies to improve performance. Drawing inspiration from molecular conformation generation methods Zhu et al. (2022), which underscores the importance of permutation invariance for symmetric atoms, we introduce a permutation loss function to enhance the robustness of the conformation prediction.
@@ 1 Introduction
Figure 1: Cases of the predicted docking results. Left: a good case with correct pocket and docking pose prediction. Right: a bad case that the predicted pocket is far away from the groud truth pocket so to a large RMSD score with incorrect ligand pose.
@@
Figure 2: Overall framework of FABind+. Left: The regression-based FABind+ consists of the original FABind design and the newly proposed dynamic pocket radius prediction and permutation loss module. Right: The sampling version of FABind+, which contains a pocket clustering module, a conformation sampling module, and a confidence selection module.
@@
As demonstrated by Corso et al. (2023a; b), a sampling model is essential for capturing multiple binding sites and conformations. Therefore, after applying the above enhancement to build a regression-based FABind+, we further exploit its sampling potential. Unlike previous sampling-based methods of training a generative model for capturing multiple conformations, we directly transform the trained regression-based FABind+ into a sampling-based model. Specifically, to address the issue of multiple binding sites, we employ a clustering method to identify all potential pocket candidates, leveraging our residue-level probability for pocket presence. Moreover, we integrate a robust yet simple sampling mechanism based on dropout Srivastava et al. (2014) to enable FABind+ to sample multiple conformations effectively. Finally, a lightweight confidence model is trained to rank the best-sampled structures.
@@
For evaluation, we conduct comprehensive experiments and analyses on the widely recognized PDBBind v2020 benchmarks Liu et al. (2017) under various settings. As a regression-based model, FABind+ outperforms all previous methods with remarkable inference speed. Moreover, by activating the sampling mode, FABind+ is capable of generating diverse and high-quality conformations, thereby further enhancing performance beyond that of the purely regression-based model.
2 Related Work
2.1 Pocket/Binding Site Prediction
Pocket prediction plays an important role in structure-based drug discovery. Early computational methods relied on hand-crafted features, using different modeling approaches such as physical chemical-based, geometric-based, and machine learning-based methods Stank et al. (2016); Weisel et al. (2007); Capra et al. (2009). For example, sequence-based techniques exploited protein sequences Taherzadeh et al. (2016); Chen & Jeong (2009), while structure-based tools examined 3D structures Le Guilloux et al. (2009); Laskowski (1995), and the integration of both sequence and 3D structures Capra et al. (2009). Recently, deep learning has achieved significant advances. Pioneering works applied convolutional neural networks (CNNs) to learn spatial patterns from protein structures VIART et al. (2020); Pu et al. (2019). DeepSite Jiménez et al. (2017) and DeepPocket Gentile et al. (2020) both apply to the sequence profiles of the protein to predict the binding hotspots. Other works includes Zhang et al. (2019); Ragoza et al. (2017); Zhang et al. (2023b). Among them, Fpocket Le Guilloux et al. (2009) and P2rank Krivák & Hoksza (2018) are two representative works that open-sourced and have been widely adopted in existing works. In this work, we carefully build a dynamic pocket prediction module and integrate this module into the entire molecular docking framework.
2.2 Molecular Docking
@@
Molecular docking, also known as protein-ligand docking, aims to predict the correct binding pose of the protein-ligand complex. Traditional methods typically use physics-informed energy functions to score, rank, and refine the ligand structures, such as AutoDock Vina Trott & Olson (2010), SMINA Koes et al. (2013), GLIDE Friesner et al. (2004), and GOLD Jones et al. (1997). Geometric deep learning has been attractive and promising for docking prediction. These methods can be divided into two categories: 1) Regression-based methods that aim to directly predict the docked ligand pose coordinates or optimize the structures with predicted pairwise distances between atoms, such as EquiBind Stärk et al. (2022), TankBind Lu et al. (2022), E3Bind Zhang et al. (2022), and KarmaDock Zhang et al. (2023a). These methods usually demonstrate clear advantages in inference speed. 2) Sampling-based methods that require multiple ligand poses sampling and then perform optimization or selection among sampled conformation candidates Nakata et al. (2023); Corso et al. (2023b); Wang et al. (2023); Guo et al. (2023). While more computationally expensive, sampling-based methods often yield more accurate predictions. FABind Pei et al. (2023) is a docking framework that is designed to be both efficient and effective. Our work follows FABind to eliminate its limitations and boost its performance. Notably, Alphafold-latest Deepmind & Labs (2023) achieves huge breakthroughs, but in close-form introduction without details. DeltaDock Yan et al. (2023) and HelixDock Liu et al. (2023) also obtain competitive results for us. However, they either generate large-scale data with simulators or add external high quality data for the training, which is not a fair comparison.
3 Method
3.1 Preliminaries
@@
Problem Definition. Let $\mathcal{G}=(\mathcal{V}:=\{\mathcal{V}^{l},\mathcal{V}^{p}\},\mathcal{E}:=\{\mathcal{E}^{l},\mathcal{E}^{p},\mathcal{E}^{lp}\})$ denote a protein-ligand complex, where $\mathcal{G}^{l}=(\mathcal{V}^{l},\mathcal{E}^{l})$ and $\mathcal{G}^{p}=(\mathcal{V}^{p},\mathcal{E}^{p})$ are ligand and protein graph, respectively. The symbols $\mathcal{V}$ and $\mathcal{E}$ represent collections of atoms (residue for protein) and bonds. $\mathcal{E}^{lp}$ in $\mathcal{E}$ is the edge collection between protein and ligand graph. Each node $v=(\mathbf{h},\mathbf{x})\in\mathcal{V}$ contains a feature vector $\mathbf{h}$ and its coordinates $\mathbf{x}\in\mathbb{R}^{3}$. For clarity, $v_{i}=(\mathbf{h}_{i},\mathbf{x}_{i})\in\mathcal{V}^{l}$ is used to denote ligand atom and $v_{j}=(\mathbf{h}_{j},\mathbf{x}_{j})\in\mathcal{V}^{p}$ for protein residue. We use $\mathcal{R}^{l}\in\mathbb{R}^{\mid\mathcal{V}^{l}\mid\times 3}$ and $\mathcal{R}^{p}\in\mathbb{R}^{\mid\mathcal{V}^{p}\mid\times 3}$ to represent the conformation of ligand and protein. Given a protein in its bound state and a flexible ligand, our objective is to learn a mapping from the random initialized ligand pose to the bounded conformation $\mathcal{R}^{l}=\{\mathbf{x}_{i}\}_{1\leq i\leq\mid\mathcal{V}^{l}\mid}$. Notably, we focus on the blind docking setting, where we possess no information regarding the binding pocket.
@@
Background of FABind. FABind Pei et al. (2023) is a novel deep learning framework that aims to provide both fast and accurate protein-ligand binding structure prediction in an end-to-end manner, which tries to overcome the low efficiency of sampling methods and also the inaccuracy of regression methods. The core innovation in FABind is its unique layer design called “FABind layer” (F). Each FABind layer $(l)$ takes in protein-ligand complex graph and protein-ligand pair embedding and outputs updated embedding and ligand structures:
$\mathbf{h}_{i}^{(l+1)},\mathbf{h}_{j}^{(l+1)},\mathbf{x}_{i}^{(l+1)},\mathbf{p}_{ij}^{(l+1)}=\text{F}(\mathbf{h}_{i}^{(l)},\mathbf{h}_{j}^{(l)},\mathbf{x}_{i}^{(l)},\mathbf{x}_{j},\mathbf{p}_{ij}^{(l)}),$
@@
where $\mathbf{p}_{ij}\in\mathbb{R}^{D\times D}$ is pair embedding for each pair of ligand and protein node and $D$ is hidden size. The FABind layer contains three key components: independent message passing, cross-attention update, and interfacial message passing. The independent message passing captures interactions within the protein and ligand separately, the cross-attention update enhances node representations by exchanging information across the protein and ligand, and the interfacial message passing focuses on modeling interactions at the protein-ligand interface.
@@
Another important technology is how FABind unifies pocket prediction and docking into a unified framework. It introduces a ligand-informed pocket prediction module that utilizes the ligand information to pinpoint the unique binding pocket. This achieves faster and more precise pocket prediction than methods relying on external pocket detectors. The predicted pocket is then leveraged for downstream docking prediction. Notably, FABind adopts a scheduled sampling Bengio et al. (2015) training strategy that gradually incorporates the predicted pocket instead of solely using the native pocket, ensuring consistency between training and inference. It also proposes distance map constraints, besides coordinate loss, to refine the predicted pose.
3.2 Enhanced FABind+

### v4

@@
Abstract.
Molecular docking is a pivotal process in drug discovery. While traditional techniques rely on extensive sampling and simulation governed by physical principles, deep learning has emerged as a promising alternative, offering improvements in both accuracy and efficiency. Building upon the foundational work of FABind, a model focused on speed and accuracy, we introduce FABind+, an enhanced iteration that significantly elevates the performance of its predecessor. We identify pocket prediction as a critical bottleneck in molecular docking and introduce an enhanced approach. In addition to the pocket prediction module, the docking module has also been upgraded with permutation loss and a more refined model design. These designs enable the regression-based FABind+ to surpass most of the generative models. In contrast, while sampling-based models often struggle with inefficiency, they excel in capturing a wide range of potential docking poses, leading to better overall performance. To bridge the gap between sampling and regression docking models, we incorporate a simple yet effective sampling technique coupled with a lightweight confidence model, transforming the regression-based FABind+ into a sampling version without requiring additional training. This involves the introduction of pocket clustering to capture multiple binding sites and dropout sampling for various conformations. The combination of a classification loss and a ranking loss enables the lightweight confidence model to select the most accurate prediction. Experimental results and analysis demonstrate that FABind+ (both the regression and sampling versions) not only significantly outperforms the original FABind, but also achieves competitive state-of-the-art performance. Our code is available at https://github.com/QizhiPei/FABind.
blind docking, molecular docking
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709253isbn: 979-8-4007-1245-6/25/08ccs: Applied computing Molecular structural biologyccs: Applied computing Bioinformaticsccs: Computing methodologies Machine learning
1. Introduction
@@
Figure 1. Pocket prediction is critical for docking. Left: a good case with correct pocket and docking pose prediction. Right: a bad case with incorrect pocket prediction.
Molecular docking is a foundational technique in drug discovery (Morris et al., 1996; Morris and Lim-Wilby, 2008; Thomsen and Christensen, 2006; AI4Science and Quantum, 2023), predicting the preferred orientation of ligands when bound to a protein target. This process is crucial for the identification and optimization of compounds with therapeutic potential. Traditional methodologies (Friesner et al., 2004; Trott and Olson, 2010; Morris et al., 1996; Verdonk et al., 2003) rely heavily on exhaustive sampling and simulation techniques based on the principles of physics and chemistry (Shoichet and Kuntz, 1993; Bernacki et al., 2005; Perez et al., 2016; Huang and Jacobson, 2007). These approaches aim to mimic the complex interactions between ligands and protein receptors to forecast optimal conformations. However, despite their extensive application, the classical methods are often criticized for their computational intensity, causing slow processing and significant resource consumption (Pagadala et al., 2017; Huang et al., 2006).
With the rise of computational power and machine learning, deep learning approaches for molecular docking have emerged recently (Zhang et al., 2023b; Lu et al., 2023; Zhou et al., 2023). These deep learning methods can be broadly categorized into two types: regression approaches and generative models. Regression approaches directly predict the coordinates (Stärk et al., 2022; Zhang et al., 2022a) or distance matrices of pocket-ligand interactions (Masters et al., 2022; Lu et al., 2022). In contrast, generative models sample multiple candidate poses and use confidence models to select the most promising ones (Corso et al., 2023b; Wang et al., 2023; Guo et al., 2023), similar to traditional methods. FABind, a notable example of a regression-based method, has demonstrated significant advancements in terms of speed and accuracy. However, despite its impressive performance, FABind’s accuracy does not fully align with the latest advancements in the field (Corso et al., 2023b; Liu et al., 2023; Yan et al., 2023).
@@
Figure 2. Overall framework of FABind+. Left: Pipeline of FABind with the newly proposed dynamic pocket radius prediction and permutation loss module. The protein-ligand complex graph, with pink and blue nodes representing protein and ligand atoms respectively, is first processed by the pocket prediction module to identify the binding pocket. The identified pocket (highlighted in dark pink) is then utilized for docking in the pose prediction module. Right: The sampling version of FABind+, which contains a pocket clustering module, a conformation sampling module, and a confidence selection module.
In this work, we introduce FABind+, which features enhanced pocket prediction capabilities and high-quality pose generation. Our initial analysis identifies pocket prediction as a key factor influencing docking accuracy, where inaccuracies can lead to suboptimal results. As illustrated in Figure 1, incorrect pocket predictions from FABind result in a significant root-mean-square deviation (RMSD), with 16.25% of poses exhibiting an RMSD greater than 10Å, largely due to these prediction errors. To mitigate this, we propose dynamically adjusting the pocket radius instead of relying on a fixed-size sphere, which allows the pocket to better encompass the entire potential ligand structure, thereby improving docking pose predictions. For the docking module, we also implement strategies to improve the performance. Inspired by previous works on molecular structure modeling (Zhu et al., 2022; Zhang et al., 2022b), which emphasize the importance of permutation invariance for symmetric atoms, we introduce a permutation loss function to increase the robustness of conformation predictions. Furthermore, we conduct model adjustments to optimize training and enhance overall performance.
@@
Demonstrated by DiffDock and subsequent works (Corso et al., 2023a, b), a sampling model is essential for capturing multiple binding sites and conformations, leading to better docking results. However, previous sampling-based methods that train generative models, mostly diffusion-based, incur high inference costs, which are undesirable in high-throughput virtual screening scenarios. Unlike previous approaches, we innovatively transform the pre-trained regression-based FABind+ into a sampling-based model without further training while maintaining the same inference time for one sampling. Specifically, to enable discovering multiple binding sites, we employ a clustering method to identify all potential pocket candidates, leveraging our residue-level probability for pockets. Besides, we integrate a simple yet effective sampling mechanism based on dropout (Srivastava et al., 2014) to enable FABind+ sample multiple conformations. Finally, a lightweight confidence model is trained on the fly to select the best-sampled structure. We additionally emphasize the importance of loss design and demonstrate that a lightweight model architecture is sufficient for docking pose selection.
@@
For evaluation, we conduct comprehensive experiments and analyses on the widely recognized PDBBind v2020 benchmarks (Liu et al., 2017) under various settings. As a regression-based model, FABind+ outperforms all previous methods with remarkable inference speed compared to its generative counterparts. Additionally, by activating the sampling mode, FABind+ is capable of generating diverse and high-quality conformations, capturing multiple pockets and multiple conformations that the regression model alone might miss, while also achieving improved docking accuracy through confidence model selection.
2. Related Work
Pocket/Binding Site Prediction. Pocket prediction plays a key role in structure-based drug discovery. Early computational methods rely on hand-crafted features and use different modeling approaches (Stank et al., 2016; Weisel et al., 2007; Capra et al., 2009). For example, sequence-based techniques exploited protein sequences (Taherzadeh et al., 2016; Chen and Jeong, 2009), while structure-based tools examined 3D structures (Le Guilloux et al., 2009; Laskowski, 1995), and the integration of both sequence and 3D structures (Capra et al., 2009). Recently, deep learning has achieved significant advances in this area. Most contemporary methods, utilizing voxel-based (VIART et al., 2020; Pu et al., 2019; Jiménez et al., 2017; Ragoza et al., 2017) and node-based representations (Le Guilloux et al., 2009; Krivák and Hoksza, 2018; Zhang et al., 2023a), implement dynamic pocket prediction, where larger pockets are predicted for larger ligands. Contrastively, Fixed pocket radius in TankBind (Lu et al., 2022), E3Bind (Zhang et al., 2022a) and FABind (Pei et al., 2023) is not sufficient. Moreover, it is crucial to model multiple binding sites within a single protein, as previous work has demonstrated (Le Guilloux et al., 2009; Krivák and Hoksza, 2018).
@@ 2. Related Work
Molecular Docking. Molecular docking predicts the correct binding pose of protein-ligand complexes. Traditional methods are typically sampling-based, which involves optimizing various initial conformations to generate different binding structures (Trott and Olson, 2010; Koes et al., 2013; Friesner et al., 2004; Jones et al., 1997). Geometric deep learning has shown promise in docking prediction, divided into two categories: (1) Regression-based methods directly predict the docked ligand pose coordinates or optimize the structures with predicted pairwise distances between atoms, such as EquiBind (Stärk et al., 2022), TankBind (Lu et al., 2022), E3Bind (Zhang et al., 2022a), and KarmaDock (Zhang et al., 2023b). These methods usually demonstrate clear advantages in inference speed. (2) Sampling-based methods require multiple ligand poses sampling and then perform optimization or selection among sampled conformation candidates (Nakata et al., 2023; Corso et al., 2023b; Wang et al., 2023; Guo et al., 2023). While more computationally expensive, sampling-based methods often yield more accurate predictions. Notably, Alphafold 3 (Abramson et al., 2024) achieves huge breakthroughs, but is not open-sourced and lacks detailed methodological transparency. DeltaDock (Yan et al., 2023) and HelixDock (Liu et al., 2023) also obtain competitive results. However, they either generate large-scale data with simulators or add extensive high-quality data for the training, which is not a fair comparison. Among related works, FABind is a regression-based approach with both efficiency and effectiveness. Our work follows FABind to enhance regression results and unlock its sampling potential.
3. Methodology
3.1. Preliminaries
@@
Problem Definition. Let $\mathcal{G}=(\mathcal{V}:=\{\mathcal{V}^{l},\mathcal{V}^{p}\},\mathcal{E}:=\{\mathcal{E}^{l},\mathcal{E}^{p},\mathcal{E}^{lp}\})$ denotes a protein-ligand complex, where $\mathcal{G}^{l}=(\mathcal{V}^{l},\mathcal{E}^{l})$ and $\mathcal{G}^{p}=(\mathcal{V}^{p},\mathcal{E}^{p})$ are ligand and protein graph, respectively. The symbols $\mathcal{V}$ and $\mathcal{E}$ represent collections of atoms (residue for protein) and bonds. $\mathcal{E}^{lp}$ in $\mathcal{E}$ is the edge collection between protein and ligand graph. Each node $v=(\mathbf{h},\mathbf{x})\in\mathcal{V}$ contains a feature vector $\mathbf{h}$ and its coordinates $\mathbf{x}\in\mathbb{R}^{3}$. For clarity, $v_{i}=(\mathbf{h}_{i},\mathbf{x}_{i})\in\mathcal{V}^{l}$ is used to denote ligand atom and $v_{j}=(\mathbf{h}_{j},\mathbf{x}_{j})\in\mathcal{V}^{p}$ for protein residue. We use $\mathcal{R}^{l}\in\mathbb{R}^{\mid\mathcal{V}^{l}\mid\times 3}$ and $\mathcal{R}^{p}\in\mathbb{R}^{\mid\mathcal{V}^{p}\mid\times 3}$ to represent the conformation of ligand and protein, respectively. Given a protein in its bound state and a flexible ligand, our objective is to learn a mapping from the randomly initialized ligand pose to the bounded conformation $\mathcal{R}^{l}=\{\mathbf{x}_{i}\}_{1\leq i\leq\mid\mathcal{V}^{l}\mid}$. Notably, we focus on the blind docking setting, where we possess no information regarding the binding pocket.
@@
Key Idea of FABind. FABind (Pei et al., 2023) is a novel deep learning framework that aims to provide both fast and accurate protein-ligand binding structure prediction in an end-to-end manner. It seeks to overcome the inefficiencies typically associated with sampling methods and the inaccuracies often seen in regression-based approaches. The first core innovation in FABind is its unique layer design called “FABind layer" (F). Each FABind layer $(l)$ processes the protein-ligand complex graph and the protein-ligand pair embedding, producing updated embeddings and ligand structures:
$\small\mathbf{h}_{i}^{(l+1)},\mathbf{h}_{j}^{(l+1)},\mathbf{x}_{i}^{(l+1)},\mathbf{p}_{ij}^{(l+1)}=\text{F}(\mathbf{h}_{i}^{(l)},\mathbf{h}_{j}^{(l)},\mathbf{x}_{i}^{(l)},\mathbf{x}_{j},\mathbf{p}_{ij}^{(l)}),$
@@
where $\mathbf{p}_{ij}\in\mathbb{R}^{D\times D}$ is the pair embedding for each pair of ligand and protein node and $D$ is the hidden size. The FABind layer contains three key components: independent message passing, cross-attention update, and interfacial message passing. The independent message passing captures interactions within the protein and ligand separately, the cross-attention update enhances node representations by exchanging information across the protein and ligand, and the interfacial message passing focuses on modeling interactions at the protein-ligand interface.
@@
Another important contribution of FABind is the decomposition of the blind docking process into pocket prediction and pocket-specific docking. The unified framework introduces a ligand-informed pocket prediction module that utilizes ligand information to identify the unique binding pocket. This approach achieves faster and more precise pocket prediction compared to methods that rely on external pocket detectors. The predicted pocket is then used for downstream docking prediction.
3.2. Enhancing FABind to FABind+